# Nyström Kernel Stein Discrepancy

### v2

@@ Abstract
Kernel methods underpin many of the most successful approaches in data science and statistics, and they allow representing probability measures as elements of a reproducing kernel Hilbert space without loss of information. Recently, the kernel Stein discrepancy (KSD), which combines Stein’s method with kernel techniques, gained considerable attention. Through the Stein operator, KSD allows the construction of powerful goodness-of-fit tests where it is sufficient to know the target distribution up to a multiplicative constant. However, the typical U- and V-statistic-based KSD estimators suffer from a quadratic runtime complexity, which hinders their application in large-scale settings. In this work, we propose a Nyström-based KSD acceleration—with runtime $\mathcal{O}\!\left(mn+m^{3}\right)$ for $n$ samples and $m\ll n$ Nyström points—, show its $\sqrt{n}$-consistency under the null with a classical sub-Gaussian assumption, and demonstrate its applicability for goodness-of-fit testing on a suite of benchmarks.
@@ 1 Introduction
The kernel mean embedding, which involves mapping probabilities into a reproducing kernel Hilbert spaces (RKHS; Aronszajn 1950) has found far-reaching applications in the last $20$ years. For example, it allows to measure the discrepancy between probability distributions through maximum mean discrepancy (MMD; Smola et al. 2007, Gretton et al. 2012), defined as the distance between the corresponding mean embeddings, which underpins powerful two-sample tests. MMD is also known as energy distance [Székely and Rizzo, 2004, 2005, Baringhaus and Franz, 2004] in the statistics literature; see Sejdinovic et al. [2013] for the equivalence. We refer to [Muandet et al., 2017] for a recent overview of kernel mean embeddings.
@@
A recent approach to tackle GoF testing involves applying a Stein operator [Stein, 1972, Chen, 2021, Anastasiou et al., 2023] to functions in an RKHS and using them as test functions to measure the discrepancy between distributions, referred to as kernel Stein discrepancies (KSD; Chwialkowski et al. 2016, Liu et al. 2016). An empirical estimator of KSD can be used as a test statistic to address the GoF problem. In particular, the Langevin Stein operator [Gorham and Mackey, 2015, Chwialkowski et al., 2016, Liu et al., 2016, Oates et al., 2017, Gorham and Mackey, 2017] in combination with the kernel mean embedding gives rise to a KSD on the Euclidean space $\mathbb{R}^{d}$, which we consider in this work. As a test statistic, KSD has many desirable properties. In particular, KSD requires only knowledge of the derivative of the score function of the target distribution — implying that KSD is agnostic to the normalization of the target and therefore does not require solving, either analytically or numerically, complex normalization integrals in Bayesian settings. This property has led to its widespread use, e.g., for assessing and improving sample quality [Gorham and Mackey, 2015, Chen et al., 2018, 2019, Futami et al., 2019, Gorham et al., 2020], validating MCMC methods [Coullon et al., 2023], comparing deep generative models [Lim et al., 2019], detecting out-of-distribution inputs [Nalisnick et al., 2019], assessing Bayesian seismic inversion [Izzatullah et al., 2020], modeling counterfactuals [Martinez-Taboada and Kennedy, 2023], and explaining predictions [Sarvmaili et al., 2024]. GoF testing with KSDs has been explored on Euclidean data [Liu et al., 2016, Chwialkowski et al., 2016], discrete data [Yang et al., 2018], point processes [Yang et al., 2019], time-to-event data [Fernandez et al., 2020], graph data [Xu and Reinert, 2021], sequential models [Baum et al., 2023], and functional data [Wynne et al., 2024]. The KSD statistic has also been extended to the conditional case [Jitkrittum et al., 2020].
@@
Estimators for Langevin Stein operator-based KSD exist. But, the classical U-statistic- [Liu et al., 2016] and V-statistic-based [Chwialkowski et al., 2016] estimators have a runtime complexity that scales quadratically with the number of samples of the sampling distribution, which limits their deployment to large-scale settings. To address this bottleneck, Chwialkowski et al. [2016] introduced a linear-time statistic that suffers from low statistical power compared to its quadratic-time counterpart. Jitkrittum et al. [2017] proposed the finite set Stein discrepancy (FSSD), a linear-time approach that replaces the RKHS-norm by the $L_{2}$-norm approximated by sampling; the sampling can either be random (FSSD-rand) or optimized w.r.t. a power proxy (FSSD-opt). Another approach [Huggins and Mackey, 2018] is employing the random Fourier feature (RFF; Rahimi and Recht 2007) method to accelerate the KSD estimation. However, it is known [Chwialkowski et al., 2015, Proposition 1] that the resulting statistic fails to distinguish a large class of measures. Huggins and Mackey [2018] generalize the idea of replacing the RKHS-norm by going from $L_{2}$-norms to $L_{p}$ ones, to obtain feature Stein discrepancies. They present an efficient approximation, random feature Stein discrepancies (RFSD), which is a (near-)linear time estimator. However, successful deployment of the method depends on a good choice of parameters, which, while the authors provide guidelines, can be challenging to select and tune in practice.
@@
Our work alleviates these severe bottlenecks. We employ the Nyström method [Williams and Seeger, 2001] to accelerate KSD estimation and show the $\sqrt{n}$-consistency of our proposed estimator under the null. The main technical challenge is that the Stein kernel (induced by the Langevin Stein operator and the original kernel) is typically unbounded while existing statistical Nyström analysis [Rudi et al., 2015, Chatalic et al., 2022, Sterge and Sriperumbudur, 2022, Kalinke and Szabó, 2023] usually considers bounded kernels. To tackle unbounded kernels, we select a classical sub-Gaussian assumption, which we impose on the feature map associated to the kernel, and show that existing methods of analysis can successfully be extended to handle this novel case. In this sense, our work, besides Della Vecchia et al. [2021], which requires a similar sub-Gaussian condition for analyzing empirical risk minimization on random subspaces, is a first step in analyzing the consistency of the unbounded case in the Nyström setting.
Our main contributions can be summarized as follows.
@@
    We prove the $\sqrt{n}$-consistency under the null of our estimator in a classical sub-Gaussian setting, which extends (in a non-trivial fashion) existing results for Nyström-based methods [Rudi et al., 2015, Chatalic et al., 2022, Sterge and Sriperumbudur, 2022, Kalinke and Szabó, 2023] focusing on bounded kernels.
@@
The paper is structured as follows. We introduce the notations used throughout the article (Section 2) followed by recalling the classical quadratic-time KSD estimators (Section 3). In Section 4.1, we detail our proposed Nyström-based estimator, alongside with its adaptation to a modified wild bootstrap goodness-of-fit test (Section 4.2), and our theoretical guarantees (Section 4.3). Experiments demonstrating the efficiency of our Nyström-KSD estimator are provided in Section 5. Proofs are deferred to the appendices.
@@ 2 Notations
Let $\left(\mathcal{X},\tau_{\mathcal{X}}\right)$ be a topological space and $\mathcal{B}\left(\tau_{\mathcal{X}}\right)$ the corresponding Borel $\sigma$-algebra. Probability measures in this article are meant w.r.t. the measurable space $\left(\mathcal{X},\mathcal{B}\left(\tau_{\mathcal{X}}\right)\right)$ and are written as $\mathcal{M}_{1}^{+}\left(\mathcal{X}\right)$; for instance, the set of Borel probability measures on $\mathbb{R}^{d}$ is $\mathcal{M}_{1}^{+}\left(\mathbb{R}^{d}\right)$. The $n$-fold product measure of $\mathbb{P}\in\mathcal{M}_{1}^{+}\left(\mathcal{X}\right)$ is denoted by $\mathbb{P}^{n}\in\mathcal{M}_{1}^{+}\left(\mathcal{X}^{n}\right)$. The product of $\mathbb{P}_{1}\in\mathcal{M}_{1}^{+}(\mathcal{X}_{1})$ and $\mathbb{P}_{2}\in\mathcal{M}_{1}^{+}(\mathcal{X}_{2})$ is written as $\mathbb{P}_{1}\otimes\mathbb{P}_{2}$ ($\in\mathcal{M}_{1}^{+}(\mathcal{X}_{1}\times\mathcal{X}_{2})$), where $\left(\mathcal{X}_{1},\tau_{\mathcal{X}_{1}}\right)$ and $\left(\mathcal{X}_{2},\tau_{\mathcal{X}_{2}}\right)$ are topological spaces. For a sequence of i.i.d. real-valued random variables $X_{n}\sim\mathbb{P}\in\mathcal{M}_{1}^{+}(\mathbb{R})$ and a sequence of positive $r_{n}$-s, $X_{n}=\mathcal{O}_{\mathbb{P}}(r_{n})$ means that $\frac{X_{n}}{r_{n}}$ is bounded in probability. The unit ball in a Hilbert space $\mathcal{H}$ is denoted by $B(\mathcal{H})=\left\{f\in\mathcal{H}\mid\left\|f\right\|_{\mathcal{H}}\leq 1\right\}$. The reproducing kernel Hilbert space with $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ as the reproducing kernel is denoted by $\mathcal{H}_{k}$. Throughout the paper, $k$ is assumed to be measurable and $\mathcal{H}_{k}$ to be separable; for instance, a continuous kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ implies the latter property [Steinwart and Christmann, 2008, Lemma 4.33]. The canonical feature map $\phi_{k}:\mathbb{R}^{d}\to\mathcal{H}_{k}$ is defined as $\mathbf{x}\mapsto k(\cdot,\mathbf{x})$; with this feature map $k(\mathbf{x},\mathbf{y})=\left\langle{k(\cdot,\mathbf{x}),k(\cdot,\mathbf{y})}\right\rangle_{\mathcal{H}_{k}}=\left\langle{\phi_{k}(\mathbf{x}),\phi_{k}(\mathbf{y})}\right\rangle_{\mathcal{H}_{k}}$ for all $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$. The Gram matrix associated with the observations $\left(\mathbf{x}_{i}\right)_{i=1}^{n}\in\left(\mathbb{R}^{d}\right)^{n}$ and kernel $k$ is $\mathbf{K}_{k,n,n}=\left[k(\mathbf{x}_{i},\mathbf{x}_{j})\right]_{i,j=1}^{n}\in\mathbb{R}^{n\times n}$. Given a closed linear subspace $U\subseteq\mathcal{H}_{k}$, the (orthogonal) projection of $h\in\mathcal{H}_{k}$ on $U$ is denoted by $P_{U}h\in U$; $u=P_{U}h$ is the unique vector such that $h-u\perp U$. For any $u\in U$, $\left\|{h-P_{U}h}\right\|_{\mathcal{H}_{k}}\leq\left\|{h-u}\right\|_{\mathcal{H}_{k}}$, that is, $P_{U}h$ is the closest element in $U$ to $h$. A linear operator $A:\mathcal{H}_{k}\to\mathcal{H}_{k}$ is called bounded if $\left\|{A}\right\|_{\mathrm{op}}:=\sup_{\left\|{h}\right\|_{\mathcal{H}_{k}}=1}\left\|{Ah}\right\|_{\mathcal{H}_{k}}<\infty$; the set of $\mathcal{H}_{k}\rightarrow\mathcal{H}_{k}$ bounded linear operators is denoted by $\mathcal{L}(\mathcal{H}_{k})$. An $A\in\mathcal{L}(\mathcal{H}_{k})$ is called positive (shortly $A\geq 0$) if it is self-adjoint ($A^{*}=A$, where $A^{*}\in\mathcal{L}(\mathcal{H}_{k})$ is defined by $\langle Af,g\rangle_{\mathcal{H}_{k}}=\langle f,A^{*}g\rangle_{\mathcal{H}_{k}}$ for all $f,g\in\mathcal{H}_{k}$), and $\langle Ah,h\rangle_{\mathcal{H}_{k}}\geq 0$ for all $h\in\mathcal{H}_{k}$. If $A\geq 0$, then there exists a unique $B\geq 0$ such that $B^{2}=A$; we write $B=A^{\frac{1}{2}}$ and call $B$ the square root of $A$. An $A\in\mathcal{L}(\mathcal{H}_{k})$ is called trace-class if $\sum_{i\in I}\langle(A^{*}A)^{\frac{1}{2}}e_{i},e_{i}\rangle_{\mathcal{H}_{k}}<\infty$ for some countable orthonormal basis (ONB) $(e_{i})_{i\in I}$ of $\mathcal{H}_{k}$, and in this case $\operatorname{tr}(A):=\sum_{i\in I}\langle Ae_{i},e_{i}\rangle_{\mathcal{H}_{k}}<\infty$. [^1] The trace-class property and the value of $\operatorname{tr}(A)$ is independent of the specific ONB chosen. The separability of $\mathcal{H}_{k}$ implies the existence of a countable ONB in it. For a self-adjoint trace-class operator $A$ with eigenvalues $(\lambda_{i})_{i\in I}$, $\operatorname{tr}(A)=\sum_{i\in I}\lambda_{i}$. An operator $A\in\mathcal{L}(\mathcal{H}_{k})$ is called compact if $\overline{\{Ah\,|\,h\in B(\mathcal{H}_{k})\}}$ is compact, where $\overline{\cdot}$ denotes the closure. A trace class operator is compact, and a compact positive operator $A$ has largest eigenvalue $\left\|{A}\right\|_{\mathrm{op}}$. For any $A\in\mathcal{L}(\mathcal{H}_{k})$, it holds that $\left\|{A^{*}A}\right\|_{\mathrm{op}}=\left\|{A}\right\|_{\mathrm{op}}^{2}$ (which is called the $C^{*}$ property).
@@
The mean embedding of a probability measure $\mathbb{P}\in\mathcal{M}_{1}^{+}(\mathbb{R}^{d})$ into the RKHS associated to kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $\mu_{k}(\mathbb{P})=\int_{\mathbb{R}^{d}}\phi_{k}(\mathbf{x})\mathrm{d}\mathbb{P}(\mathbf{x})\in\mathcal{H}_{k}$, where the integral is meant in Bochner’s sense [Diestel and Uhl, 1977, Chapter II.2]. The mean element $\mu_{k}(\mathbb{P})$ exists iff. $\int_{\mathbb{R}^{d}}\left\|{\phi_{k}(\mathbf{x})}\right\|_{\mathcal{H}_{k}}\mathrm{d}\mathbb{P}(\mathbf{x})<\infty$ [Diestel and Uhl, 1977, p. 45; Theorem 2]; this condition is satisfied for instance when $\phi_{k}$ is bounded, in other words $\sup_{\mathbf{x}\in\mathbb{R}^{d}}\left\|{\phi_{k}(\mathbf{x})}\right\|_{\mathcal{H}_{k}}<\infty$.

### v4

@@ Abstract
Kernel methods underpin many of the most successful approaches in data science and statistics, and they allow representing probability measures as elements of a reproducing kernel Hilbert space without loss of information. Recently, the kernel Stein discrepancy (KSD), which combines Stein’s method with the flexibility of kernel techniques, gained considerable attention. Through the Stein operator, KSD allows the construction of powerful goodness-of-fit tests where it is sufficient to know the target distribution up to a multiplicative constant. However, the typical U- and V-statistic-based KSD estimators suffer from a quadratic runtime complexity, which hinders their application in large-scale settings. In this work, we propose a Nyström-based KSD acceleration—with runtime $\mathcal{O}\!\left(mn+m^{3}\right)$ for $n$ samples and $m\ll n$ Nyström points—, show its $\sqrt{n}$-consistency with a classical sub-Gaussian assumption, and demonstrate its applicability for goodness-of-fit testing on a suite of benchmarks. We also show the $\sqrt{n}$-consistency of the quadratic-time KSD estimator.
@@ 1 INTRODUCTION
The kernel mean embedding, which involves mapping probability distributions into a reproducing kernel Hilbert space (RKHS; Aronszajn 1950) has found various far-reaching applications in the last $20$ years. For example, it allows to measure the discrepancy between probability distributions through maximum mean discrepancy (MMD; Smola et al. 2007, Gretton et al. 2012), defined as the distance between the corresponding mean embeddings, which underpins powerful two-sample tests. MMD is also known as energy distance [Székely and Rizzo, 2004, 2005, Baringhaus and Franz, 2004] in the statistics literature; see Sejdinovic et al. [2013] for the equivalence. We refer to [Muandet et al., 2017] for a recent overview of kernel mean embeddings.
@@
A recent approach to tackle GoF testing involves applying a Stein operator [Stein, 1972, Chen, 2021, Anastasiou et al., 2023] to functions in an RKHS and using them as test functions to measure the discrepancy between distributions, referred to as kernel Stein discrepancies (KSD; Chwialkowski et al. 2016, Liu et al. 2016). An empirical estimator of KSD can be used as a test statistic to address the GoF problem. In particular, the Langevin Stein operator [Gorham and Mackey, 2015, Chwialkowski et al., 2016, Liu et al., 2016, Oates et al., 2017, Gorham and Mackey, 2017] in combination with the kernel mean embedding gives rise to a KSD on the Euclidean space $\mathbb{R}^{d}$, which we consider in this work. As a test statistic, KSD has many desirable properties. In particular, KSD requires only knowledge of the derivative of the score function of the target distribution — implying that KSD is agnostic to the normalization of the target and therefore does not require solving, either analytically or numerically, complex normalization integrals in Bayesian settings. This property has led to its widespread use, e.g., for assessing and improving sample quality [Gorham and Mackey, 2015, Chen et al., 2018, 2019, Futami et al., 2019, Gorham et al., 2020], validating MCMC methods [Coullon et al., 2023], comparing deep generative models [Lim et al., 2019], detecting out-of-distribution inputs [Nalisnick et al., 2019], assessing Bayesian seismic inversion [Izzatullah et al., 2020], modeling counterfactuals [Martinez-Taboada and Kennedy, 2024], and explaining predictions [Sarvmaili et al., 2025]. GoF testing with KSDs has been explored on Euclidean data [Liu et al., 2016, Chwialkowski et al., 2016], discrete data [Yang et al., 2018], point processes [Yang et al., 2019], time-to-event data [Fernandez et al., 2020], graph data [Xu and Reinert, 2021], sequential models [Baum et al., 2023], and functional data [Wynne et al., 2024]. The KSD statistic has also been extended to the conditional case [Jitkrittum et al., 2020].
@@
Estimators for Langevin Stein operator-based KSD exist. But, the classical U-statistic- [Liu et al., 2016] and V-statistic-based [Chwialkowski et al., 2016] estimators have a runtime complexity that scales quadratically with the number of samples of the sampling distribution, which limits their deployment to large-scale settings. To address this bottleneck, Chwialkowski et al. [2016] introduced a linear-time statistic that suffers from low statistical power compared to its quadratic-time counterpart. Jitkrittum et al. [2017b] proposed the finite set Stein discrepancy (FSSD), a linear-time approach that replaces the RKHS-norm by the $L_{2}$-norm approximated by sampling; the sampling can either be random (FSSD-rand) or optimized w.r.t. a power proxy (FSSD-opt). Another approach [Huggins and Mackey, 2018] is employing the random Fourier feature (RFF; Rahimi and Recht 2007, Sriperumbudur and Szabó 2015) method to accelerate the KSD estimation. However, it is known [Chwialkowski et al., 2015, Proposition 1] that the resulting statistic fails to distinguish a large class of measures. Huggins and Mackey [2018] generalize the idea of replacing the RKHS-norm by going from $L_{2}$-norms to $L_{p}$ ones, to obtain feature Stein discrepancies. They present an efficient approximation, random feature Stein discrepancies (RFSD), which is a (near-)linear time estimator. However, successful deployment of the method depends on a good choice of parameters, which, while the authors provide guidelines, can be challenging to select and tune in practice.
@@
Our work alleviates these severe bottlenecks. We employ the Nyström method [Williams and Seeger, 2001] to accelerate KSD estimation and show the $\sqrt{n}$-consistency of our proposed estimator. The main technical challenge is that the Stein kernel (induced by the Langevin Stein operator and the original kernel) is typically unbounded while existing statistical Nyström analysis [Rudi et al., 2015, Chatalic et al., 2022, Sterge and Sriperumbudur, 2022, Kalinke and Szabó, 2023, Chatalic et al., 2025] usually considers bounded kernels. To tackle unbounded kernels, we select a classical sub-Gaussian assumption, which we impose on the feature map associated to the kernel, and show that existing methods of analysis can successfully be extended to handle this novel case. In this sense, our work, besides Della Vecchia et al. [2021], which requires a similar sub-Gaussian condition for analyzing empirical risk minimization on random subspaces, is a first step in analyzing the consistency of the unbounded case in the Nyström setting.
Our main contributions are the following.
@@
    We prove the $\sqrt{n}$-consistency of our estimator in a classical sub-Gaussian setting, which extends (in a non-trivial fashion) existing results for Nyström-based methods [Rudi et al., 2015, Chatalic et al., 2022, Sterge and Sriperumbudur, 2022, Kalinke and Szabó, 2023] focusing on bounded kernels.
@@
The paper is structured as follows. We introduce the notations used throughout the article (Section 2) followed by recalling the classical quadratic-time KSD estimators (Section 3). In Section 4.1, we detail our proposed Nyström-based estimator, alongside with its adaptation to a modified wild bootstrap goodness-of-fit test (Section 4.2), and our theoretical guarantees (Section 4.3). Experiments demonstrating the efficiency of our Nyström-KSD estimator are provided in Section 5. Limitations are in Section 6. Proofs and additional experiments are deferred to the appendices.
@@ 2 NOTATIONS
Let $\left(\mathcal{X},\tau_{\mathcal{X}}\right)$ be a topological space and $\mathcal{B}\left(\tau_{\mathcal{X}}\right)$ the corresponding Borel $\sigma$-algebra. Probability measures considered in this article are meant w.r.t. the measurable space $\left(\mathcal{X},\mathcal{B}\left(\tau_{\mathcal{X}}\right)\right)$ and are written as $\mathcal{M}_{1}^{+}\left(\mathcal{X}\right)$; for instance, the set of Borel probability measures on $\mathbb{R}^{d}$ is $\mathcal{M}_{1}^{+}\left(\mathbb{R}^{d}\right)$. The $n$-fold product measure of $\mathbb{P}\in\mathcal{M}_{1}^{+}\left(\mathcal{X}\right)$ is denoted by $\mathbb{P}^{n}\in\mathcal{M}_{1}^{+}\left(\mathcal{X}^{n}\right)$. The product of $\mathbb{P}_{1}\in\mathcal{M}_{1}^{+}(\mathcal{X}_{1})$ and $\mathbb{P}_{2}\in\mathcal{M}_{1}^{+}(\mathcal{X}_{2})$ is written as $\mathbb{P}_{1}\otimes\mathbb{P}_{2}$ ($\in\mathcal{M}_{1}^{+}(\mathcal{X}_{1}\times\mathcal{X}_{2})$), where $\left(\mathcal{X}_{1},\tau_{\mathcal{X}_{1}}\right)$ and $\left(\mathcal{X}_{2},\tau_{\mathcal{X}_{2}}\right)$ are topological spaces. For a sequence of i.i.d. real-valued random variables $X_{n}\sim\mathbb{P}\in\mathcal{M}_{1}^{+}(\mathbb{R})$ and a sequence of positive $r_{n}$-s, $X_{n}=\mathcal{O}_{\mathbb{P}}(r_{n})$ means that $\frac{X_{n}}{r_{n}}$ is bounded in probability. The unit ball in a Hilbert space $\mathcal{H}$ is denoted by $B(\mathcal{H})=\left\{f\in\mathcal{H}\mid\left\|f\right\|_{\mathcal{H}}\leq 1\right\}$. The reproducing kernel Hilbert space with $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ as the reproducing kernel is denoted by $\mathcal{H}_{k}$. Throughout the paper, $k$ is assumed to be measurable and $\mathcal{H}_{k}$ to be separable. [^1] For instance, a continuous kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ implies both properties; see Steinwart and Christmann [2008, Lemma 4.33] for separability. Given a closed linear subspace $U\subseteq\mathcal{H}_{k}$, the (orthogonal) projection of $h\in\mathcal{H}_{k}$ on $U$ is denoted by $P_{U}h\in U$; $u=P_{U}h$ is the unique vector such that $h-u\perp U$. For any $u\in U$, $\left\|{h-P_{U}h}\right\|_{\mathcal{H}_{k}}\leq\left\|{h-u}\right\|_{\mathcal{H}_{k}}$, that is, $P_{U}h$ is the closest element in $U$ to $h$. A linear operator $A:\mathcal{H}_{k}\to\mathcal{H}_{k}$ is called bounded if $\left\|{A}\right\|_{\mathrm{op}}:=\sup_{\left\|{h}\right\|_{\mathcal{H}_{k}}=1}\left\|{Ah}\right\|_{\mathcal{H}_{k}}<\infty$; the set of $\mathcal{H}_{k}\rightarrow\mathcal{H}_{k}$ bounded linear operators is denoted by $\mathcal{L}(\mathcal{H}_{k})$. An $A\in\mathcal{L}(\mathcal{H}_{k})$ is called positive (shortly $A\geq 0$) if it is self-adjoint ($A^{*}=A$, where $A^{*}\in\mathcal{L}(\mathcal{H}_{k})$ is defined by $\langle Af,g\rangle_{\mathcal{H}_{k}}=\langle f,A^{*}g\rangle_{\mathcal{H}_{k}}$ for all $f,g\in\mathcal{H}_{k}$), and $\langle Ah,h\rangle_{\mathcal{H}_{k}}\geq 0$ for all $h\in\mathcal{H}_{k}$. If $A\geq 0$, then there exists a unique $B\geq 0$ such that $B^{2}=A$; we write $B=A^{\frac{1}{2}}$ and call $B$ the square root of $A$. An $A\in\mathcal{L}(\mathcal{H}_{k})$ is called trace-class if $\sum_{i\in I}\langle(A^{*}A)^{\frac{1}{2}}e_{i},e_{i}\rangle_{\mathcal{H}_{k}}<\infty$ for some countable orthonormal basis (ONB) $(e_{i})_{i\in I}$ of $\mathcal{H}_{k}$, and in this case $\operatorname{tr}(A):=\sum_{i\in I}\langle Ae_{i},e_{i}\rangle_{\mathcal{H}_{k}}<\infty$. [^2] The trace-class property and the value of $\operatorname{tr}(A)$ is independent of the specific ONB chosen. The separability of $\mathcal{H}_{k}$ implies the existence of a countable ONB in it. For a self-adjoint trace-class operator $A$ with eigenvalues $(\lambda_{i})_{i\in I}$, $\operatorname{tr}(A)=\sum_{i\in I}\lambda_{i}$. An operator $A\in\mathcal{L}(\mathcal{H}_{k})$ is called compact if $\overline{\{Ah\,|\,h\in B(\mathcal{H}_{k})\}}$ is compact, where $\overline{\cdot}$ denotes the closure. A trace class operator is compact, and a compact positive operator $A$ has largest eigenvalue $\left\|{A}\right\|_{\mathrm{op}}$. For any $A\in\mathcal{L}(\mathcal{H}_{k})$, it holds that $\left\|{A^{*}A}\right\|_{\mathrm{op}}=\left\|{A}\right\|_{\mathrm{op}}^{2}$ (which is called the $C^{*}$ property).
@@
The mean embedding of a probability measure $\mathbb{P}\in\mathcal{M}_{1}^{+}(\mathbb{R}^{d})$ into the RKHS associated to kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $\mu_{k}(\mathbb{P})=\int_{\mathbb{R}^{d}}k\left(\cdot,\mathbf{x}\right)\mathrm{d}\mathbb{P}(\mathbf{x})\in\mathcal{H}_{k}$, where the integral is meant in Bochner’s sense [Diestel and Uhl, 1977, Chapter II.2]. The mean element $\mu_{k}(\mathbb{P})$ exists iff. $\int_{\mathbb{R}^{d}}\left\|{k\left(\cdot,\mathbf{x}\right)}\right\|_{\mathcal{H}_{k}}\mathrm{d}\mathbb{P}(\mathbf{x})<\infty$ [Diestel and Uhl, 1977, p. 45; Theorem 2].